OpenStudy (aryavegapunk):

Calculate work needed to take an iphone 5 from the surface of the Earth to infinity. Google needed constants and parameters.

OpenStudy (anonymous):

Work is \[ \int F~dx \]

OpenStudy (anonymous):

We can use \[ F = G\frac{Mm}{x^2} \]

OpenStudy (anonymous):

So I think we will do: \[ W = \int_{x_{\text{Earth radius}}}^{\infty}G\frac{Mm}{x^2}dx \]

OpenStudy (anonymous):

So I think we will do: \[ W = GMm\int_{x_{\text{Earth radius}}}^{\infty}\frac{dx}{x^2}=GMm \left[-\frac{1}{x}\right|_{x_{\text{Earth radius}}}^{\infty} = \frac{GMm}{x_{\text{Earth radius}}} \]

OpenStudy (anonymous):

\(G\) is just a gravitational constant, \(M\) and \(m\) are the masses of Earth and the iPhone 5